40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed May 25 16:56:44 2022, Paco, Configuration, BHD, IFO recovery - IMC alignment 
    Reply  Wed May 25 17:34:47 2022, yuta, Configuration, BHD, IFO recovery - IFO alignment Screenshot_2022-05-25_17-47-57.png
       Reply  Thu May 26 19:55:43 2022, yuta, Configuration, BHD, Oplevs centered, BHD DCPDs are now online elog_1Y2.JPGelog_BHD.JPGelog_box.JPGScreenshot_2022-05-26_17-37-27_IFOaligned_OplevCentered.pngScreenshot_2022-05-26_20-35-02.png
          Reply  Fri May 27 17:45:53 2022, yuta, Configuration, BHD, BHD camera installed, GRY aligned 
             Reply  Thu Jun 2 20:05:37 2022, yuta, Configuration, PSL, IMC input power recovered to 1W, some alignment works Before.JPGAfter.JPG
                Reply  Fri Jun 3 12:13:58 2022, Paco, Configuration, CDS, Fix RFM channels  SoGreen.png
                   Reply  Fri Jun 3 15:22:51 2022, yuta, Update, LSC, Both arms locked with POY/POX, IR beam centered on TMs with ASS IRBeamsOnTMs.JPGScreenshot_2022-06-03_15-03-51.png
                      Reply  Fri Jun 3 17:42:50 2022, yuta, Update, LSC, MICH locks with AS55_Q Screenshot_2022-06-03_17-41-55.png
                         Reply  Sun Jun 5 19:46:40 2022, Paco, Update, LSC, Fixed IMC Trans sum issue 
                         Reply  Mon Jun 6 13:35:11 2022, Paco, Update, LSC, First calibrated spectra of MICH at AS55 Q Screenshot_2022-06-06_17-30-16_MICHOLTF.pngCalibrated_MICH_ERR.pdf
                            Reply  Fri Jun 17 16:22:21 2022, yuta, Update, LSC, Actuator calibration of BS. ITMX, ITMY, updated MICH displacement spectra from c1cal 6x
          Reply  Thu Jun 16 18:23:15 2022, Paco, Configuration, BHD, Recovering LO beam in BHD DCPDs Screenshot_2022-06-16_18-29-14_BHDLObeamISBACK.png
          Reply  Tue Jun 21 14:17:50 2022, yuta, Configuration, BHD, BHD DCPDs re-routed to c1sus2 C1X07ADC1.JPGBHDDCPDs.JPG
Message ID: 16880     Entry time: Fri May 27 17:45:53 2022     In reply to: 16877     Reply to this: 16886
Author: yuta 
Type: Configuration 
Category: BHD 
Subject: BHD camera installed, GRY aligned 

[JC, Paco, Yuta]

After the IFO recovery (elog 40m/16881), we installed an analog camera for BHD fringe using a BNC cable for old SRMF camera so that we can see it from the control room.
We also aligned AS-LO using LO1,LO2 and AS4.
We then aligned GRY injection to get maximum GTRY.

Maximum TEM00s right now are
 C1:SUS-ETMX_TRX_OUT_DQ ~0.1
 C1:SUS-ETMY_TRY_OUT_DQ ~0.05
 C1:ALS-TRX_OUT_DQ ~0.20
 C1:ALS-TRY_OUT_DQ ~0.18

ELOG V3.1.3-